3(-9n + 5) + -4 = -16
Answer by jessica43(140)

The first step to solve this is to use the distributive property. The Distributive Property is helps you get rid of parentheses:
a(b + c) = ab + ac
So in this problem:
3(-9n + 5) + -4 = -16
3(-9n) + (3)(5) + (-4) = -16
-27n + 15 - 4 = -16
-27n + 11 = -16 (now subtract 11 from each side)
-27n = -27 (now divide each side by -27)
n = 1
So the answer is n = 1.